OpenStudy (anonymous):

PLEASE HELP!!! Solve logx + 5 36 = 2.

OpenStudy (nurali):

let "log[b](x)" = "log, base b, of x" so if you mean: log[x + 5](36) = 2 then 36 = (x + 5)^2 ← basic definition 6 = x + 5 1 = x